# Near-linear Time Dispersion of Mobile Agents

### v2

@@ Abstract
Consider that there are $k\leq n$ agents in a simple, connected, and undirected graph $G=(V,E)$ with $n$ nodes and $m$ edges. The goal of the dispersion problem is to move these $k$ agents to distinct nodes. Agents can communicate only when they are at the same node, and no other means of communication such as whiteboards are available. We assume that the agents operate synchronously. We consider two scenarios: when all agents are initially located at any single node (rooted setting) and when they are initially distributed over any one or more nodes (general setting). Kshemkalyani and Sharma presented a dispersion algorithm for the general setting, which uses $O(m_{k})$ time and $\log(k+\delta)$ bits of memory per agent [OPODIS 2021]. Here, $m_{k}$ is the maximum number of edges in any induced subgraph of $G$ with $k$ nodes, and $\delta$ is the maximum degree of $G$. This algorithm is the fastest in the literature, as no algorithm with $o(m_{k})$ time has been discovered even for the rooted setting. In this paper, we present faster algorithms for both the rooted and general settings. First, we present an algorithm for the rooted setting that solves the dispersion problem in $O(k\log\min(k,\delta))=O(k\log k)$ time using $O(\log\delta)$ bits of memory per agent. Next, we propose an algorithm for the general setting that achieves dispersion in $O(k(\log k)\cdot(\log\min(k,\delta))=O(k\log^{2}k)$ time using $O(\log(k+\delta))$ bits. Finally, for the rooted setting, we give a time-optimal, i.e., $O(k)$-time, algorithm with $O(\delta)$ bits of space per agent.
@@ 1 Introduction
In this paper, we focus on the dispersion problem involving mobile robots, which we refer to as mobile agents, or simply, agents. At the start of an execution, $k$ agents are arbitrarily positioned within an undirected graph. The objective of this problem is to ensure that all agents are located at distinct nodes. This problem was originally proposed by Augustine and Moses Jr. [2] in 2018. A particularly intriguing aspect of this problem is the unique nature of the computation model. In contrast to many other models involving mobile agents on graphs, we do not have access to the identifiers of the nodes, nor can we utilize a local memory at each node. In this setting, an agent cannot retrieve or store any information from or on a node when it visits. However, each of the $k$ agents possesses a unique identifier and can communicate with each other when they stay at the same node in a graph. The agents are tasked with collaboratively solving a common problem through direct communication with each other.
Table 1: Dispersion of mobile agents on an arbitrary graph ($\tau=\min(k,\delta)$)
Memory per agent Time General/Rooted Knowledge Async./Sync.
[2] $O(\log\delta)$ $O(m_{k})$ rooted - async.
[7] $O(d+\delta\log k)$ $O(d\delta(d+\delta))$ rooted - async.
Theorem 1 $O(\log\delta)$ $O(k\log\tau)$ rooted - sync.
Theorem 3 $O(\delta)$ $O(k)$ rooted - sync.
[2] $O(\log n)$ $O(\delta^{d})$ general - async.
[2] $O(k\log(k+\delta))$ $O(m_{k})$ general - async.
[5] $O(d\log\delta)$ $O(\delta^{d})$ general - async.
[5] $O(\log(k+\delta))$ $O(m_{k}\cdot k)$ general - async.
[6] $O(\log(k+\delta))$ $O(m_{k}\log k)$ general $m_{k},k$ sync.
[14] $O(\log(k+\delta))$ $O(m_{k}\log k)$ general - sync.
[9] $O(\log(k+\delta))$ $O(m_{k})$ general - async.
Theorem 2 $O(\log(k+\delta))$ $O(k\log^{2}k)$ general - sync.
Lower bound any $\Omega(k)$ any any any
@@
Several algorithms have been introduced in the literature to solve the dispersion problem. We define a parameter $m_{k}$ as the maximum number of edges among any $k$ nodes in a graph. This parameter serves as an upper bound on the number of edges between two nodes, each accommodating at least one agent. Consequently, $m_{k}$ frequently appears in the time complexities of several algorithms. This is because (i) solving the dispersion problem essentially requires finding $k$ distinct nodes, and (ii) the simple depth-first search (DFS), employed as a submodule by many dispersion algorithms, needs to explore $m_{k}$ edges to find $k$ nodes. The dispersion problem has been examined in two different contexts within the literature: the rooted setting and the general setting. In the rooted setting, all $k$ agents initially reside at a single node. On the other hand, the general setting imposes no restrictions on the initial placement of the $k$ agents.
@@
Table 1 provides a summary of various dispersion algorithms found in the literature, all designed for arbitrary graphs.111 In this paper, including Table 1, we evaluate space complexity excluding the memory space each agent needs to store its own identifier. If you want to consider that space, you can add an extra $O(\log k)$ bits per agent to the listed space complexity. Augustine and Moses Jr. [2] introduced a simple algorithm, based on depth-first search (DFS), for the rooted setting. This algorithm solves the dispersion problem in $O(m_{k})$ time using $O(\log\delta)$ bits of space per agent, where $\delta$ is the maximum degree of a graph. They then extended this algorithm to the general setting. This expanded version also achieves dispersion in $O(m_{k})$ time, but it requires roughly $k$ times more space, i.e., $O(k\log(k+\delta))$ bits. Kshemkalyani and Ali [5] offered an algorithm that accomplishes dispersion in the general setting using less space: $O(\log(k+\delta))$ bits per agent. However, it requires more time, i.e., $O(m_{k}\cdot k)$. These two algorithms for the general setting offer a trade-off between time and space. The first is faster but needs more space, while the second is slower but more memory-efficient. Kshemkalyani, Molla, and Sharma [6] found a middle ground with an algorithm that runs in $O(m_{k}\log k)$ time and uses $O(\log(k+\delta))$ bits of each agent’s memory. This algorithm, however, requires global knowledge, such as asymptotically tight upper bounds on $m_{k}$ and $k$, to attain its time upper bound. Shintaku, Sudo, Kakugawa, and Masuzawa [14] managed to eliminate this requirement for global knowledge. More recently, Kshemkalyani and Sharma [9] removed the $\log k$ factor from the running time. This algorithm also works in an asynchronous setting, meaning the agents do not need to share a common clock. We show a naive lower bound on the running time by considering a path: any dispersion algorithm requires at least $\Omega(k)$ time. Thus, there is still a significant gap between the best known upper bound $O(m_{k})$ and this lower bound of $\Omega(k)$ because $m_{k}=\Theta(k^{2})$ holds in many graph classes. Note that $m_{k}=\Theta(k^{2})$ may hold in a sparse graph when $k=O(\sqrt{n})$, where $n$ is the number of nodes in the graph.
@@ 1.1 Our Contribution
In this paper, we drastically reduce the gap between the upper bound $O(m_{k})$ and the lower bound $\Omega(k)$ mentioned earlier. Let $\tau=\min(k,\delta)$. We present two algorithms: one for the rooted setting that achieves dispersion in $O(k\log\tau)=O(k\log k)$ time using $O(\log\delta)$ bits, and the other for the general setting that achieves dispersion in $O(k(\log k)\cdot(\log\tau))=O(k\log^{2}k)$ time using $O(\log(k+\delta))$ bits. The upper bounds obtained here match the lower bounds in both the rooted and general settings when ignoring poly-logarithmic factors.
@@
It might seem that the overhead can be eliminated by using the DFS parallelization method proposed by Kshemkalyani and Sharma[9], instead of the method of Shintaku et al [14]. However, this is not the case because our HEO-DFS is not compatible with the parallelization method of Kshemkalyani and Sharma. Specifically, their method entails a process such that one DFS absorbs another when multiple DFSs collide. During this process, it is necessary to gather the agents in the absorbed side to a single node, which requires $\Theta(m_{k})$ time. Our speed-up idea effectively works for finding an unvisited neighbor, but it does not work for the acceleration of gathering agents dispersed on multiple nodes. Therefore, it is unlikely that our HEO-DFS can be combined with the method of Kshemkalyani and Sharma.
@@
While the two algorithms mentioned above are nearly time-optimal, i.e., requiring $O(k\cdot\log^{c}k)$ time for some constant $c$, they are not entirely optimal. We also demonstrate that in the rooted setting, a time-optimal algorithm based on the HEO-DFS can be achieved if significantly more space is available, specifically $O(\delta)$ bits per agent.
1.2 Related Work
@@
The dispersion problem has been studied not only for arbitrary undirected graphs but also for graphs with restricted topologies such as trees [2], grids [6, 8], and dynamic rings [1]. Several studies have explored randomized algorithms to minimize the space complexity of dispersion [11, 4]. There have also been several studies on fault-tolerant dispersion [10, 3]. Additionally, Kshemkalyani et al. [7] introduced the global communication model, in which all agents can communicate with each other regardless of their current locations.
@@
Exploration by a mobile agent is closely related to the dispersion problem. The exploration problem requires an agent to visit all nodes of a graph. Many studies have addressed the exploration problem, and numerous efficient algorithms, both in terms of time and space, have been presented in the literature [13, 12, 15]. In contrast to exploration, the dispersion problem only requires finding $k$ nodes, and we can use $k$ agents to achieve this. Our HEO-DFS take advantage of these differences to solve the dispersion problem efficiently.
@@ 2 Preliminaries
We evaluate the time complexity or running time of algorithm $\mathcal{A}$ as the maximum number of steps until $\Xi_{\mathcal{A}}(C_{0})$ reaches a legitimate configuration, where the maximum is taken over all initial configurations $C_{0}$. Let $\mathcal{M}^{\prime}_{\mathcal{A}}\subseteq\mathcal{M}_{\mathcal{A}}$ be the set of all agent-states that can appear in any possible execution of $\mathcal{A}$ starting from any initial configuration. We evaluate the space complexity or memory space of algorithm $\mathcal{A}$ as $\log_{2}|\mathcal{M}^{\prime}_{\mathcal{A}}|$, i.e., the maximum number of bits required to represent an agent-state that may appear in those executions. This implies that we exclude the size of the working memory used for deciding the destination and updating states, and the size of the space for storing input information. In particular, the space complexity can be $o(\log k)$ or $o(\log\delta)$, while each agent reads its identifier and the incoming port at each time step.222 We adopt this definition following the convention in the study of mobile agents. Some papers dealing with the dispersion problem count the number of bits required for storing the agent identifier and the incoming port.
@@
For any dispersion algorithm $\mathcal{A}$, there exists a graph $G$ such that an execution of $\mathcal{A}$ requires $\Omega(k)$ time steps to achieve dispersion on both the rooted and general settings.
In the two algorithms we present in this paper, $\mathbf{RootedDispersion}$ and $\mathbf{GeneralDispersion}$, each agent maintains a variable $a.\mathtt{settled}\in\{\bot,\top\}$. We say that an agent $a$ is a settler when $a.\mathtt{settled}=\top$, and an explorer otherwise. All agents are explorers initially. Once an explorer becomes a settler, it never becomes an explorer again. Let $t$ be the time at which an agent $a$ becomes a settler. Thereafter, we call the location of $a$ at that time, i.e.$\nu(a,t)$, the home of $a$. Formally, $a$’s home at $t^{\prime}\geq 0$, denoted by $\xi(a,t^{\prime})$, is defined as $\xi(a,t^{\prime})=\bot$ if $t^{\prime}<t$ and $\xi(a,t^{\prime})=\nu(a,t)$ otherwise. It is worth mentioning that a settler may temporarily leave its home. Hence $\xi(a,t^{\prime})=\nu(a,t^{\prime})$ may not always hold even after $a$ becomes a settler, i.e., even if $t^{\prime}\geq t$. However, by definition, no agent changes its home. We say that an agent $a$ settles when it becomes a settler.

### v3

@@ Abstract
Consider that there are $k\leq n$ agents in a simple, connected, and undirected graph $G=(V,E)$ with $n$ nodes and $m$ edges. The goal of the dispersion problem is to move these $k$ agents to mutually distinct nodes. Agents can communicate only when they are at the same node, and no other communication means, such as whiteboards, are available. We assume that the agents operate synchronously. We consider two scenarios: when all agents are initially located at a single node (rooted setting) and when they are initially distributed over one or more nodes (general setting). Kshemkalyani and Sharma presented a dispersion algorithm for the general setting, which uses $O(m_{k})$ time and $\log(k+\Delta)$ bits of memory per agent [OPODIS 2021], where $m_{k}$ is the maximum number of edges in any induced subgraph of $G$ with $k$ nodes, and $\Delta$ is the maximum degree of $G$. This algorithm is currently the fastest in the literature, as no $o(m_{k})$-time algorithm has been discovered, even for the rooted setting. In this paper, we present significantly faster algorithms for both the rooted and the general settings. First, we present an algorithm for the rooted setting that solves the dispersion problem in $O(k\log\min(k,\Delta))=O(k\log k)$ time using $O(\log(k+\Delta))$ bits of memory per agent. Next, we propose an algorithm for the general setting that achieves dispersion in $O(k\log k\cdot\log\min(k,\Delta))=O(k\log^{2}k)$ time using $O(\log(k+\Delta))$ bits. Finally, for the rooted setting, we give a time-optimal (i.e., $O(k)$-time) algorithm with $O(\Delta+\log k)$ bits of space per agent. All algorithms presented in this paper work only in the synchronous setting, while several algorithms in the literature, including the one given by Kshemkalyani and Sharma at OPODIS 2021, work in the asynchronous setting.
@@ 1 Introduction
In this paper, we focus on the dispersion problem involving mobile entities, referred to as mobile agents, or simply, agents. At the start of an execution, $k$ agents are arbitrarily positioned at nodes of an undirected graph $G=(V,E)$ with $n$ nodes and $m$ edges. The objective is to ensure that all agents are located at mutually distinct nodes. This problem was originally proposed by Augustine and Moses Jr. [2] in 2018. A particularly intriguing aspect of this problem is the unique computation model. Unlike many other models involving mobile agents on graphs, we do not have access to node identifiers, nor can we use local memory at each node. In this setting, an agent cannot retrieve or store any information from or on a node when it visits. However, each of the $k$ agents possesses a unique identifier and can communicate with each other when they are at the same node in the graph. The agents must collaboratively solve the dispersion problem through this direct communication.
Table 1: Dispersion of mobile agents on an arbitrary graph ($\tau=\min(k,\Delta)$) The algorithm of [7] needs to know an asymptotically tight upper bound on $m_{k}$ in advance. (Since $m_{k}\leq\min(m,k\Delta,\binom{k}{2})$, knowing tight upper bounds on $m$, $k$, and $\Delta$ is sufficient, but it increases the running time to $O(\min(m,k\Delta,\binom{k}{2})\log k)$.) The $\log k$ term can be eliminated from the space complexities marked with daggers ($\dagger$) if you choose to disregard the memory space required for each agent to store its own identifier. For example, the proposed algorithm mentioned in Theorem 1 requires only $O(\log\Delta)$ bits per agent. The space complexity of the first algorithm given by [6], which is marked with a double dagger ($\ddagger$), can be decreased to $O(k\log\Delta)$ if we assume that the number of possible agent-identifiers is $O(k)$. All algorithms listed in this table are deterministic.
Memory per agent Time General/Rooted Async./Sync.
[2] $O(\log(k+\Delta))$dagger $O(m_{k})$ rooted async.
[10] $O(D+\Delta\log k)$ $O(D\Delta(D+\Delta))$ rooted async.
Theorem 1 $O(\log(k+\Delta))$dagger $O(k\log\tau)$ rooted sync.
Theorem 3 $O(\Delta+\log k)$dagger $O(k)$ rooted sync.
[6] $O(k\log(k+\Delta))$ddagger $O(m_{k})$ general async.
[6] $O(D\log\Delta+\log k)$dagger $O(\Delta^{D})$ general async.
[6] $O(\log(k+\Delta))$ $O(m_{k}\cdot k)$ general async.
[7] $O(\log(k+\Delta))$ $O(m_{k}\log k)$ general sync.
[16] $O(\log(k+\Delta))$ $O(m_{k}\log k)$ general sync.
[11] $O(\log(k+\Delta))$ $O(m_{k})$ general async.
Theorem 2 $O(\log(k+\Delta))$ $O(k\log^{2}k)$ general sync.
Lower bound any $\Omega(k)$ any any
@@
Several algorithms have been introduced in the literature to solve the dispersion problem. This problem has been examined in two different contexts within the literature: the rooted setting and the general setting. In the rooted setting, all $k$ agents initially reside at a single node. On the other hand, the general setting imposes no restrictions on the initial placement of the $k$ agents. For any $i\in[1,n]$, we define $m_{i}$ as the maximum number of edges in any $i$-node induced subgraph of $G$. The parameter $m_{k}$, where $k$ is the number of agents, serves as an upper bound on the number of edges connecting two nodes, each hosting at least one agent, in any configuration. Consequently, $m_{k}$ frequently appears in the time complexities of dispersion algorithms. This is because (i) solving the dispersion problem essentially requires finding $k$ distinct nodes, and (ii) the simple depth-first search (DFS), employed as a submodule by many dispersion algorithms, needs to explore $m_{k}$ edges to find $k$ nodes.
@@
Table 1 provides a summary of various dispersion algorithms found in the literature, all designed for arbitrary graphs. Here, $\Delta$ and $D$ are the maximum degree and the diameter of a graph, respectively, and $\tau=\min(k,\Delta)$. Augustine and Moses Jr. [2] introduced a simple algorithm, based on depth-first search (DFS), for the rooted setting. This algorithm solves the dispersion problem in $O(m_{k})$ time using $O(\log\Delta)$ bits of space per agent. Kshemkalyani and Ali [6] provides two algorithms that accomplish dispersion in the general setting: an $O(m_{k})$-time and $O(k\log(k+\Delta))$-space algorithm, and an $O(m_{k}\cdot k)$-time and $O(\log(k+\Delta))$-space algorithm, offering a trade-off between time and space. The first is faster but needs more space, while the second is slower but more memory-efficient. Kshemkalyani, Molla, and Sharma [7] found a middle ground with an algorithm that runs in $O(m_{k}\log k)$ time and uses $O(\log(k+\Delta))$ bits of each agent’s memory. This algorithm, however, requires a priori global knowledge, asymptotically tight upper bounds on $m_{k}$, to attain its time upper bound. Shintaku, Sudo, Kakugawa, and Masuzawa [16] managed to eliminate this requirement for global knowledge. More recently, Kshemkalyani and Sharma [11] removed the $\log k$ factor from the running time. This algorithm also works in an asynchronous setting, meaning the agents do not need to share a common clock. Any dispersion algorithm requires at least $\Omega(k)$ time, which is almost trivial, but we will provide a proof for completeness in this paper. No other lower bounds on the time complexity of dispersion have been established in the literature. Thus, there is still a significant gap between the best known upper bound $O(m_{k})$ and this lower bound of $\Omega(k)$ because $m_{k}=\Theta(k^{2})$ holds in many graph classes. Note that $m_{k}=\Theta(k^{2})$ may hold even in a sparse graph when $k=O(\sqrt{n})$.
@@
Note (Space Complexity).
Conforming to the convention in the studies of mobile agents [5], this paper, including Table 1, evaluates the space complexity of an algorithm as the maximum size of persistent memory needed by an agent during its execution. Persistent memory refers to the information an agent carries when it moves from one node to another and does not include the working memory used for local computations at nodes. This persistent memory includes the space required to store its own identifier. Since the $k$ agents are labeled with unique identifiers, every algorithm requires $O(\log k)$ bits per agent.
Note (Parameter $m_{k}$).
The parameter $m_{k}$ is introduced in this paper and has not been previously utilized in the literature. Traditionally, the running times of DFS-based algorithms are represented using the parameter $\min(m,k\Delta)$ or $\min(m,k\Delta,\binom{k}{2})$, which are always greater than or equal to $m_{k}$. The parameter $m_{k}$ may be better to represent them because there are some graph classes where $m_{k}=o(k^{2})$ while $\min(m,k\Delta,\binom{k}{2})=\Omega(k^{2})$.
@@ 1.1 Our Contribution
In this paper, we drastically reduce the gap between the upper bound $O(m_{k})$ and the lower bound $\Omega(k)$ mentioned earlier. As previously noted, $m_{k}=\Theta(k^{2})$ may hold even in sparse graphs, making this gap significant. Let $\tau=\min(k,\Delta)$. We present two algorithms: one for the rooted setting that achieves dispersion in $O(k\log\tau)=O(k\log k)$ time using $O(\log\Delta)$ bits, and the other for the general setting that achieves dispersion in $O(k\log k\cdot\log\tau)=O(k\log^{2}k)$ time using $O(\log(k+\Delta))$ bits. The upper bounds obtained here match the lower bounds in both the rooted and the general settings when ignoring poly-logarithmic factors.
@@
It might seem that the overhead can be eliminated by using the DFS parallelization method proposed by Kshemkalyani and Sharma [11], instead of the method of Shintaku et al. [16]. However, this is not the case because our HEO-DFS is not compatible with the parallelization method of Kshemkalyani and Sharma. Specifically, their method entails a process such that one DFS absorbs another when multiple DFSs collide. During this process, it is necessary to gather the agents in the absorbed side to a single node, which requires $\Theta(m_{k})$ time. Our speed-up idea effectively works for finding an unvisited neighbor, but it does not work for the acceleration of gathering agents dispersed on multiple nodes. Therefore, it is unlikely that our HEO-DFS can be combined with the method of Kshemkalyani and Sharma.
@@
The two algorithms mentioned above are nearly time-optimal, i.e., requiring $O(k\cdot\log^{c}k)$ time for some constant $c$. We also demonstrate that in the rooted setting, a time-optimal algorithm based on the HEO-DFS can be achieved if significantly more space is available, specifically $O(\Delta)$ bits per agent.
To the best of our knowledge, HEO-DFS is a novel approach, and no similar techniques have been used in the literature concerning mobile agents and mobile robots. While we demonstrate that HEO-DFS significantly reduces the running time of dispersion algorithms, this technique may also prove useful for addressing other fundamental problems such as exploration and gathering.
A drawback of our HEO-DFS is that it fundamentally requires a synchronous model, even in the rooted setting; i.e., it does not function in an asynchronous model. In HEO-DFS, we attempt to find an unvisited neighbor of the current location with the help of agents settled on other neighbors. These agents must return to their homes once the probing process is completed. In an asynchronous model, unsettled agents (and/or helping agents) may visit those homes before their owners return, disrupting the consistent behavior of HEO-DFS. Therefore, the algorithm provided by Kshemkalyani and Sharma [11] remains the fastest for the asynchronous model. It is still an open question whether there exists a $o(k^{2})$-time algorithm that accommodates asynchronicity.
Note (Termination).
In this paper, we do not explicitly mention how the agents terminate the execution of a given algorithm. In many cases, termination is straightforward without any additional assumptions in the rooted setting, while in the general setting, additional assumptions are required. Specifically, in the general setting, all algorithms listed in Table 1, except for the $O(\Delta^{d})$-time algorithm presented in [8] 111 However, in this algorithm, the agents do not terminate simultaneously, and they require the ability to detect whether or not there is a terminated agent at the current location. , require both a synchronous setting and global knowledge such as (asymptotically tight upper bounds on) $m_{k}$ and $k$. With these assumptions, the agents can easily terminate simultaneously after a sufficiently large number of steps, e.g.$\Theta(k\log^{2}k)$ steps in our algorithm for the general setting. Thus, when termination is required, our general setting algorithm no longer exhibits disadvantages compared to existing algorithms: all existing algorithms, except for the $O(\Delta^{D})$-time one [8], also require a synchronous setting (and some global knowledge).
For completeness, we present how the agents terminate in our algorithms for the rooted setting, which is almost trivial, in Section A in the appendix.
1.2 Further Related Work
@@
The dispersion problem has been studied not only for arbitrary undirected graphs but also for graphs with restricted topologies such as trees [2], grids [7, 9], and dynamic rings [1]. Additionally, several studies have explored randomized algorithms to minimize the space complexity of dispersion [13, 4], and others have focused on fault-tolerant dispersion [12, 3]. Kshemkalyani et al. [10] introduced the global communication model, where all agents can communicate with each other regardless of their locations. In contrast, the standard model, where only the agents co-located at the same node can communicate with each other, is sometimes referred to as the local communication model. All algorithms listed in Table 1 assume the local communication model and are deterministic.
@@
Exploration by a single mobile agent is closely related to the dispersion problem. The exploration problem requires an agent to visit all nodes of a graph. Many studies have addressed the exploration problem, and numerous efficient algorithms, both in terms of time and space, have been presented in the literature [15, 14, 17, 18]. In contrast to exploration, the dispersion problem only requires finding $k$ nodes, and we can use $k$ agents to achieve this. Our HEO-DFS take advantage of these differences to solve the dispersion problem efficiently.
1.3 Organization of this paper
The remainder of this paper is structured as follows:
  • Section 2 introduces the basic definitions, the model of computation, and the formal problem specification.
  • Sections 3 and 4 present near time-optimal dispersion algorithms in the rooted and general settings, respectively, utilizing $O(\log(k+\Delta))$ bits per agent.
  • Section 5 describes a time-optimal algorithm for the rooted setting, achieved by increasing memory usage to $O(\Delta+\log k)$ bits per agent, trading off increased memory for optimal time performance.
  • Section 6 summarizes the contributions of this work and discusses the conditions under which the HEO-DFS technique can accommodate the lack of synchronicity.
  • The Appendix details the termination procedures for the two proposed algorithms designed for the rooted setting.
@@ 2 Preliminaries
We evaluate the time complexity or running time of algorithm $\mathcal{A}$ as the maximum number of steps until $\Xi_{\mathcal{A}}(C_{0})$ reaches a legitimate configuration, where the maximum is taken over all initial configurations $C_{0}$. Let $\mathcal{M}^{\prime}_{\mathcal{A}}\subseteq\mathcal{M}_{\mathcal{A}}$ be the set of all agent-states that can appear in any possible execution of $\mathcal{A}$ starting from any initial configuration. We evaluate the space complexity or memory space of algorithm $\mathcal{A}$ as $\log_{2}|\mathcal{M}^{\prime}_{\mathcal{A}}|+\log_{2}\mathrm{id}_{\mathrm{max}}$, i.e., the maximum number of bits required to represent an agent-state that may appear in those executions, plus the number of bits required for each agent to store its own identifier. This implies that we exclude the size of the working memory used for deciding the destination and updating states, as well as the space for storing input information, except for the agent’s own identifier.
@@
For any dispersion algorithm $\mathcal{A}$, there exists a graph $G$ such that an execution of $\mathcal{A}$ requires $\Omega(k)$ time steps to achieve dispersion on both the rooted and the general settings.
In the two algorithms we present in this paper, $\mathbf{RootedDisp}$ and $\mathbf{GeneralDisp}$, each agent maintains a variable $a.\mathtt{settled}\in\{\bot,\top\}$. We say that an agent $a$ is a settler when $a.\mathtt{settled}=\top$, and an explorer otherwise. All agents are explorers initially. Once an explorer becomes a settler, it never becomes an explorer again. Let $t$ be the time at which an agent $a$ becomes a settler. Thereafter, we call the location of $a$ at that time, i.e.$\nu(a,t)$, the home of $a$. Formally, $a$’s home at $t^{\prime}\geq 0$, denoted by $\xi(a,t^{\prime})$, is defined as $\xi(a,t^{\prime})=\bot$ if $t^{\prime}<t$ and $\xi(a,t^{\prime})=\nu(a,t)$ otherwise. It is worth mentioning that a settler may temporarily leave its home. Hence $\xi(a,t^{\prime})=\nu(a,t^{\prime})$ may not always hold even after $a$ becomes a settler, i.e., even if $t^{\prime}\geq t$. However, by definition, no agent changes its home. We say that an agent $a$ settles when it becomes a settler.